# R u s s i a n   A c a d e m y   o f   S c i e n c e s  S i b e r i a n    B r a n c h

### v37

@@ 2.19. Are finitely generated subgroups of a free soluble group finitely separable?
No (J. S. Wilson, Arch. Math., 25 (1974), 574–577).
@@
Yes, there exists, for any prime $p$  (V. D. Mazurov, Ural Gos. Univ. Mat. Zap., 7, no. 3 (1969/70), 129–132 (Russian)).
@@
No. For example, the group $G=\left<a_{n},\,b_{n}\;(n\in\mathbb{Z}),\;c,\,d\mid[a_{n},b_{n}]=c,\;\;a_{n}^{d}=a_{n+1},\right$. $\left.b_{n}^{d}=b_{n+1},\;\;[a_{n},a_{m}]=[b_{n},b_{m}]=[a_{n},c]=[b_{n},c]=1\;\;(n,m\in\mathbb{Z})\right$> is soluble and orderable with finitely many convex subgroups; but it is not residually finite and hence is not linear over a field. (V. A. Churkin, 1969.)
@@ 2.39. Does there exist a non-finitely-axiomatizable variety of
a) Yes, there exists (S. I. Adian, Math. USSR–Izv., 4 (1970), 721–739;  A. Yu. Ol’shanskiĭ, Math. USSR–Izv., 4 (1970), 381–389).
b) Yes, there exists (A. Ya. Belov, Fundam. Prikl. Mat., 5, no. 1 (1999), 47–66 (Russian),   Sb. Math., 191, no. 3–4 (2000), 329–340;  A. V. Grishin, Fundam. Prikl. Mat., 5, no. 1 (1999), 101–118 (Russian);   V. V. Shchigolev, Fundam. Prikl. Mat., 5, no. 1 (1999), 307–312 (Russian)). But every variety of associative algebras over a field of characteristic 0 has a finite basis of identities (A. R. Kemer, Algebra and Logic, 26, no. 5 (1987), 362–397).
c) Yes, there exists (M. R. Vaughan-Lee, Quart. J. Math., 21 (1970), 297–308).
@@ 2.41. Is the variety generated by
Yes, there exists; for example, ${\mathfrak{M}}={\mathfrak{A}}_{p}{\mathfrak{A}}_{p}$, where ${\mathfrak{A}}_{p}$ is the variety of all abelian groups of prime exponent $p$ (Yu. M. Gorchakov, Talk of June 21, 1967, Krasnoyarsk).
@@ 2.47. In which abelian groups is the lattice of all fully invariant subgroups a chain?
Yes, there exists (G. M. Bergman, Pacific J. Math., 147, no. 2 (1991), 243–248).
@@ 2.79. Do there exist divisible (simple) groups with maximal subgroups?
Yes, there exist, for example, (V. G. Sokolov, Algebra and Logic, 7 (1968), 122–126).
@@ G. Baumslag
Yes, it is; for infinite cyclic extension this is proved in (M. Feighn, M. Handel, Ann. Math., 149, no. 3 (1999), 1061–1077), while another easier argument applies for finite cyclic extension.
@@ 4.54. Are any two minimal relation-modules of a finite group isomorphic?
a) Yes, there exists (A. Yu. Ol’shanskiĭ, Math. USSR–Sb., 54 (1986), 57–80).
b) Yes, there exists (P. A. Kozhevnikov, On varieties of groups of large odd exponent, Dep. 1612-V00, VINITI, Moscow, 2000 (Russian);  S. V. Ivanov, A. M. Storozhev, Contemp. Math., 360 (2004), 55–62).
@@
Yes, it exists (A. Yu. Ol’shanskii, M. V. Sapir, The conjugacy problem and Higman embeddings (Memoirs AMS, 804), 2004, 133 p.).
@@
Yes, there exists (Yu. G. Kleiman, Trans. Moscow Math. Soc., 1983, no. 2, 63–110).
@@ 5.51. Does a non-abelian free group have an independent basis for its quasi-laws?
Yes, there exists. Every non-soluble variety $\mathfrak{V}$} contains a non-trivial group coinciding with the derived subgroup, the direct limit of the spectrum $F\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}F\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}\ldots$, where $F$ is the free group in $\mathfrak{V}$} on the free generators $x_{1},x_{2},\ldots$ and $\varphi$ is the homomorphism given by $x_{i}\rightarrow[x_{2i-1},x_{2i}]$. As shown in (Yu. P. Razmyslov, Algebra and Logic, 10 (1971), 21–29) the Kostrikin variety of locally nilpotent groups of prime exponent $p\geqslant 5$ is unsoluble. (E. I. Khukhro, I. V. L’vov, Letter of June, 19, 1976.) The same was also proved in (Yu. A. Kolmakov, Math. Notes, 35, no. 5–6 (1984), 389–391). An example answering the question was also produced in (M. R. Vaughan-Lee, J. Wiegold, Bull. London Math. Soc., 13, no. 1 (1981), 45–46).
@@
Yes, there exists (B. Stellmacher, Israel J. Math., 94 (1996), 367–379).
@@
Yes, they exists (N. S. Romanovskiĭ, Sibirsk. Mat. Zh., 26, no. 4 (1985), 194–195 (Russian)).
@@
6.38. a) Let $k$ be a (commutative) field. Find all irreducible subgroups $G$ of $GL_{n}(k)$ having the property that $G\cap C\neq\varnothing$ for every conjugacy class $C$ of $GL_{n}(k)$. I conjecture that $G=GL_{n}(k)$ except in case $n={\rm char}\,k=2$, the field $k$ is quadratically closed, and $G$ is conjugate to the group of all matrices of the form $\left(\!\!\begin{array}[]{cc}\alpha&0\cr 0&\beta\end{array}\!\!\right)$, $\left(\!\!\begin{array}[]{cc}0&\alpha\cr\beta&0\end{array}\!\!\right)$ where $\alpha\neq 0$ and $\beta\neq 0$. Peter M. Neumann
@@
No, it must not (V. N. Obraztsov, Proc. Roy. Soc. Edinburgh, 123 (1993), 839–855).
@@ 6.58. Are
Yes, they exist, in both cases (A. Yu. Ol’shanskiĭ, The geometry of defining relations in groups, Kluwer, Dordrecht, 1991).
@@ 7.4. Is a finitely generated group with quadratic growth almost abelian?
Yes, it exists (A. Yu. Ol’shanskii, M. V. Sapir, Publ. Math. Inst. Hautes Étud. Sci., 96 (2002), 43–169).
@@
Yes, it exists (V. S. Guba, Math. USSR–Izv., 29 (1986), 233–277).
@@
Yes, they exist (M. I. Anokhin, Moscow Univ. Math. Bull., 51, no. 1 (1996), 48–49).
@@ 9.8. Does there exist a finitely generated simple group of intermediate growth?
Yes, they exist (D. Osin, Ann. Math., 172, no. 1 (2010), 1–39).
@@
Yes, they exist (J. Avenhaus, K. Madlener, F. Otto, Trans. Amer. Math. Soc., 297 (1986), 427–443).
@@ 9.58. Can a product of non-local formations of finite groups be local?
Yes, there exists (V. N. Obraztsov, Math. USSR–Sb., 66 (1989), 541–553).
@@ 10.23. Is it true that extraction of roots in braid groups is unique up to conjugation?
a) Yes, it does (O. Bogopolskii, A. Martino, O. Maslakova, E. Ventura, Bull. London Math. Soc., 38, no. 5 (2006), 787–794); further generalizations and applications are in (O. Bogopolski, A. Martino, E. Ventura, Trans. Amer. Math. Soc., 362 (2010), 2003–2036).
@@ 10.28. Is it true that finite strongly regular graphs with $\lambda=1$ have rank 3?
Yes, it exists (P. A. Linnell, J. Algebra, 248 (2002) 605–607).
@@ 10.56. Is the lattice of formations of finite nilpotent groups of class $\leqslant 4$ distributive?
No, not always (O. V. Bogopol’skiĭ, Geom. Dedicata, 94 (2002), 63–89).
@@
Yes, there exists (N. T. Vorob’ev, A. N. Skiba, Problems in Algebra, 8, Gomel’, 1995, 55–58 (Russian)).
@@
Yes, it exists (A. Minasyan, Comment. Math. Helv., 84 (2009), 259–296).
@@
b) Yes, it exists (I. Malinowska, Rend. Sem. Mat. Padova, 91 (1994), 265–271).
@@
11.70. b) Let $F$ be an infinite field or a skew-field. What conditions on the subring $R$ of $F$ will ensure that PGL(d+1,R)PGL(d+1,\,R) is flag-transitive on the projective $d$-space $PG(d,F)$? Peter M. Neumann, C. E. Praeger
@@
Yes, it exists (V. A. Sereda, A. I. Sozutov, Algebra and Logic, 45, no. 2 (2006), 134–138).
11.103. Is a 2-group satisfying the minimum condition for centralizers necessarily locally finite? J. Wilson
@@
Yes, they exist (N. Yu. Makarenko, E. I. Khukhro, Algebra and Logic, 45 (2006), 326–343).
@@
No, it may contain elements of order 22  (O. Kharlampovich, A. Myasnikov, in: Knots, braids, and mapping class groups — papers dedicated to Joan S. Birman (New York, 1998), Amer. Math. Soc., Providence, RI, 2001, 77–83).
@@ 13.66. Let $F$ be a
b) Yes, it exists (P.-E. Caprace, K. Fujiwara, Geom. Funct. Anal., 19 (2010), 1296–1319).
@@
Yes, it exists: the Nottingham group over ${\mathbb{F}}_{p}$ for $p>2$, which was shown to be finitely presented in (M. V. Ershov, J. London Math. Soc., 71 (2005), 362–378). (M. Ershov, Letter of 19.10.2009.)
@@
Yes, it exists (M. Kassabov, Invent. Math., 170 (2007), 297–326;  M. Ershov, A. Jaikin-Zapirain, Invent. Math., 179, 303–347).
@@ A. Moretó
15.35. Let $F$ be the free group of finite rank $r$ with basis $\{x_{1},\ldots,x_{r}\}$. Is it true that there exists a number $C=C(r)$ such that any reduced word of length $n>1$ in the $x_{i}$ lies outside some subgroup of $F$ of index at most $C\log n$? O. V. Bogopol’skiĭ
@@ 16.35. Is every finitely presented soluble group nilpotent-by-nilpotent-by-finite?
No: for a prime $p$ the group $G=\langle a,b,c,d\mid b^{a}=b^{p},\;\,c^{a}=c^{p},\;\,d^{a}=d,\;\,c^{b}=c,\break d^{c}=dd^{b},\;\,[d,d^{b}]=1,\;\,d^{p}=1\rangle$ is soluble of derived length 3, but is not metanilpotent-by-finite. (V. V. Bludov, Letter of 27 April 2007.)
@@ 16.52. Is every finitely presented elementary amenable group solvable-by-finite?
Yes, it exists (V. V. Bludov, J. Group Theory, 12, no. 4 (2009), 579–590).
@@ 16.71. Is the elementary theory of a torsion-free hyperbolic group decidable?
Yes, it can: in (D. J. Collins, Proc. London Math. Soc. (3), 36 (1978), no. 3, 480–493) it was proved that for integers $r,s$ such that $(r,s)=\nobreak 1$, $|r|\neq|s|$, and rsr-s is even, the group $G=\langle a,b\mid a^{-1}b^{r}a=b^{s}\rangle$ is isomorphic to ${\rm Aut}({\rm Aut}(G))$ (V. A. Churkin, Letter of 5 April 2006).
@@
Yes, there exist (R. I. Grigorchuk, Algebra Discrete Math., 2009, no. 4 (2009), 78–96 (subm. 25 November 2009);  A. Minasyan, A. Yu. Olshanskii, D. Sonkin, Groups Geom. Dynam., 3, no. 3 (2009) 423–452 (subm. 21 April 2008)).
@@
Yes, there exist: for example, $G_{1}=L_{15}(2^{60}).3$ and $G_{2}=L_{15}(2^{60}).5$ (M. A. Grechkoseeva, Algebra and Logic, 47, no. 4 (2008), 229–241).

### v38

@@ 2.19. Are finitely generated subgroups of a free soluble group finitely separable?
No, they do not (J. S. Wilson, Arch. Math., 25 (1974), 574–577).
@@
Yes, there does, for any prime $p$ (V. D. Mazurov, Ural Gos. Univ. Mat. Zap., 7, no. 3 (1969/70), 129–132 (Russian)).
@@
No, not every. For example, the group $G=\left<a_{n},\,b_{n}\;(n\in\mathbb{Z}),\;c,\,d\mid[a_{n},b_{n}]=c,\right$. $\left.a_{n}^{d}=a_{n+1},\;\;b_{n}^{d}=b_{n+1},\;\;[a_{n},a_{m}]=[b_{n},b_{m}]=[a_{n},c]=[b_{n},c]=1\;\;(n,m\in\mathbb{Z})\right$> is soluble and orderable with finitely many convex subgroups, but it is not residually finite and hence is not linear over a field. (V. A. Churkin, 1969.)
@@ 2.39. Does there exist a non-finitely-axiomatizable variety of
a) Yes, there does (S. I. Adian, Math. USSR–Izv., 4 (1970), 721–739; A. Yu. Ol’shanskiĭ, Math. USSR–Izv., 4 (1970), 381–389).
b) Yes, there does (A. Ya. Belov, Fundam. Prikl. Mat., 5, no. 1 (1999), 47–66 (Russian), Sb. Math., 191, no. 3–4 (2000), 329–340; A. V. Grishin, Fundam. Prikl. Mat., 5, no. 1 (1999), 101–118 (Russian); V. V. Shchigolev, Fundam. Prikl. Mat., 5, no. 1 (1999), 307–312 (Russian)). But every variety of associative algebras over a field of characteristic 0 has a finite basis of identities (A. R. Kemer, Algebra and Logic, 26, no. 5 (1987), 362–397).
c) Yes, there does (M. R. Vaughan-Lee, Quart. J. Math., 21 (1970), 297–308).
@@ 2.41. Is the variety generated by
Yes, there does; for example, ${\mathfrak{M}}={\mathfrak{A}}_{p}{\mathfrak{A}}_{p}$, where ${\mathfrak{A}}_{p}$ is the variety of all abelian groups of prime exponent $p$ (Yu. M. Gorchakov, Talk of June 21, 1967, Krasnoyarsk).
@@ 2.47. In which abelian groups is the lattice of all fully invariant subgroups a chain?
Yes, there does (G. M. Bergman, Pacific J. Math., 147, no. 2 (1991), 243–248).
@@ 2.79. Do there exist divisible (simple) groups with maximal subgroups?
Yes, there do (V. G. Sokolov, Algebra Logic, 7 (1968), 122–126).
@@ G. Baumslag
Yes, it is; moreover, every finitely generated subgroup is finitely presented. For infinite cyclic extensions this is proved in (M. Feighn, M. Handel, Ann. Math., 149, no. 3 (1999), 1061–1077), while another easier argument applies for finite cyclic extensions.
@@ 4.54. Are any two minimal relation-modules of a finite group isomorphic?
a) Yes, there does (A. Yu. Ol’shanskiĭ, Math. USSR–Sb., 54 (1986), 57–80).
b) Yes, there does (P. A. Kozhevnikov, On varieties of groups of large odd exponent, Dep. 1612-V00, VINITI, Moscow, 2000 (Russian); S. V. Ivanov, A. M. Storozhev, Contemp. Math., 360 (2004), 55–62).
@@
Yes, there does (A. Yu. Ol’shanskii, M. V. Sapir, The conjugacy problem and Higman embeddings (Memoirs AMS, 804), 2004, 133 p.).
@@
Yes, there does (Yu. G. Kleiman, Trans. Moscow Math. Soc., 1983, no. 2, 63–110).
@@ 5.51. Does a non-abelian free group have an independent basis for its quasi-laws?
Yes, there does. Every non-soluble variety $\mathfrak{V}$} contains a non-trivial group coinciding with the derived subgroup, the direct limit of the spectrum $F\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}F\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}\ldots$, where $F$ is the free group in $\mathfrak{V}$} on the free generators $x_{1},x_{2},\ldots$ and $\varphi$ is the homomorphism given by $x_{i}\rightarrow[x_{2i-1},x_{2i}]$. As shown in (Yu. P. Razmyslov, Algebra and Logic, 10 (1971), 21–29) the Kostrikin variety of locally nilpotent groups of prime exponent $p\geqslant 5$ is unsoluble. (E. I. Khukhro, I. V. L’vov, Letter of June, 19, 1976.) The same was also proved in (Yu. A. Kolmakov, Math. Notes, 35, no. 5–6 (1984), 389–391). An example answering the question was also produced in (M. R. Vaughan-Lee, J. Wiegold, Bull. London Math. Soc., 13, no. 1 (1981), 45–46).
@@
Yes, there does (B. Stellmacher, Israel J. Math., 94 (1996), 367–379).
@@
Yes, such $\gamma$-diagrams exist (N. S. Romanovskiĭ, Sibirsk. Mat. Zh., 26, no. 4 (1985), 194–195 (Russian)).
@@
6.38. a) Let $k$ be a (commutative) field. Find all irreducible subgroups $G$ of $GL_{n}(k)$ having the property that $G\cap C\neq\varnothing$ for every conjugacy class $C$ of $GL_{n}(k)$. I conjecture that $G=GL_{n}(k)$ except in case $n={\rm char}\,k=2$, the field $k$ is quadratically closed, and $G$ is conjugate to the group of all matrices of the form $\left(\!\!\begin{array}[]{cc}\alpha&0\cr 0&\beta\end{array}\!\!\right)$, $\left(\!\!\begin{array}[]{cc}0&\alpha\cr\beta&0\end{array}\!\!\right)$ where $\alpha\neq 0$ and $\beta\neq 0$. P. M. Neumann
@@
No, it need not (V. N. Obraztsov, Proc. Roy. Soc. Edinburgh, 123 (1993), 839–855).
@@ 6.58. Are
Yes, such monsters exist, in both cases (A. Yu. Ol’shanskiĭ, The geometry of defining relations in groups, Kluwer, Dordrecht, 1991).
@@ 7.4. Is a finitely generated group with quadratic growth almost abelian?
Yes, there does (A. Yu. Ol’shanskii, M. V. Sapir, Publ. Math. Inst. Hautes Étud. Sci., 96 (2002), 43–169).
@@
Yes, there does (V. S. Guba, Math. USSR–Izv., 29 (1986), 233–277).
@@
Yes, there do (M. I. Anokhin, Moscow Univ. Math. Bull., 51, no. 1 (1996), 48–49).
@@ 9.8. Does there exist a finitely generated simple group of intermediate growth?
Yes, there do (D. Osin, Ann. Math., 172, no. 1 (2010), 1–39).
@@
Yes, there do (J. Avenhaus, K. Madlener, F. Otto, Trans. Amer. Math. Soc., 297 (1986), 427–443).
@@ 9.58. Can a product of non-local formations of finite groups be local?
Yes, there does (V. N. Obraztsov, Math. USSR–Sb., 66 (1989), 541–553).
@@ 10.23. Is it true that extraction of roots in braid groups is unique up to conjugation?
a) Yes, it does (O. Bogopolski, A. Martino, O. Maslakova, E. Ventura, Bull. London Math. Soc., 38, no. 5 (2006), 787–794); further generalizations and applications are in (O. Bogopolski, A. Martino, E. Ventura, Trans. Amer. Math. Soc., 362 (2010), 2003–2036).
@@ 10.28. Is it true that finite strongly regular graphs with $\lambda=1$ have rank 3?
Yes, there does (P. A. Linnell, J. Algebra, 248 (2002) 605–607).
@@ 10.56. Is the lattice of formations of finite nilpotent groups of class $\leqslant 4$ distributive?
No, not always (O. V. Bogopolski, Geom. Dedicata, 94 (2002), 63–89).
@@
Yes, there does (N. T. Vorob’ev, A. N. Skiba, Problems in Algebra, 8, Gomel’, 1995, 55–58 (Russian)).
@@
Yes, there does (A. Minasyan, Comment. Math. Helv., 84 (2009), 259–296).
@@
b) Yes, there does (I. Malinowska, Rend. Sem. Mat. Padova, 91 (1994), 265–271).
@@
11.70. b) Let $F$ be an infinite field or a skew-field. What conditions on the subring $R$ of $F$ will ensure that PGL(d+1,R)PGL(d+1,\,R) is flag-transitive on the projective $d$-space $PG(d,F)$? P. M. Neumann, C. E. Praeger
@@
Yes, there is a Golod group with trivial centre (V. A. Sereda, A. I. Sozutov, Algebra and Logic, 45, no. 2 (2006), 134–138).
11.103. Is a 2-group satisfying the minimum condition for centralizers necessarily locally finite? J. S. Wilson
@@
Yes, there do (N. Yu. Makarenko, E. I. Khukhro, Algebra and Logic, 45 (2006), 326–343).
@@
No, it may contain elements of order $2$ (O. Kharlampovich, A. Myasnikov, in: Knots, braids, and mapping class groups — papers dedicated to Joan S. Birman (New York, 1998), Amer. Math. Soc., Providence, RI, 2001, 77–83).
@@ 13.66. Let $F$ be a
b) Yes, there does (P.-E. Caprace, K. Fujiwara, Geom. Funct. Anal., 19 (2010), 1296–1319).
@@
Yes, there does: the Nottingham group over ${\mathbb{F}}_{p}$ for $p>2$, which was shown to be finitely presented in (M. V. Ershov, J. London Math. Soc., 71 (2005), 362–378). (M. Ershov, Letter of 19.10.2009.)
@@
Yes, there does (M. Kassabov, Invent. Math., 170 (2007), 297–326; M. Ershov, A. Jaikin-Zapirain, Invent. Math., 179, 303–347).
@@ A. Moretó
15.35. Let $F$ be the free group of finite rank $r$ with basis $\{x_{1},\ldots,x_{r}\}$. Is it true that there exists a number $C=C(r)$ such that any reduced word of length $n>1$ in the $x_{i}$ lies outside some subgroup of $F$ of index at most $C\log n$? O. V. Bogopolski
@@ 16.35. Is every finitely presented soluble group nilpotent-by-nilpotent-by-finite?
No: for a prime $p$ the group $G=\langle a,b,c,d\mid b^{a}=b^{p},\;\,c^{a}=c^{p},\;\,d^{a}=d,\;\,c^{b}=c,\penalty-10000d^{c}=dd^{b},\;\,[d,d^{b}]=1,\;\,d^{p}=1\rangle$ is soluble of derived length 3, but is not metanilpotent-by-finite. (V. V. Bludov, Letter of 27 April 2007.)
@@ 16.52. Is every finitely presented elementary amenable group solvable-by-finite?
Yes, there does (V. V. Bludov, J. Group Theory, 12, no. 4 (2009), 579–590).
@@ 16.71. Is the elementary theory of a torsion-free hyperbolic group decidable?
[…]
@@
Yes, there do (R. I. Grigorchuk, Algebra Discrete Math., 2009, no. 4 (2009), 78–96 (subm. 25 November 2009); A. Minasyan, A. Yu. Olshanskii, D. Sonkin, Groups Geom. Dynam., 3, no. 3 (2009) 423–452 (subm. 21 April 2008)).
@@
Yes, there do: for example, $G_{1}=L_{15}(2^{60}).3$ and $G_{2}=L_{15}(2^{60}).5$ (M. A. Grechkoseeva, Algebra and Logic, 47, no. 4 (2008), 229–241).